OpenStudy (anonymous):

Let X,the number of flaws on a surface of the randomly selected boiler of a certain type,have a Possion distribution with parameter u=5,use Cumulative Poisson Probabilities table to compute the probablity of: P(5

OpenStudy (anonymous):

How can I use Statistical table for it?

OpenStudy (amistre64):

you got a pic of the table?

OpenStudy (anonymous):

I don't know which table I should use,it's a book

OpenStudy (amistre64):

https://onlinecourses.science.psu.edu/stat414/node/82 this might be useful for some light reading

OpenStudy (anonymous):

Thanks a lot

OpenStudy (amistre64):

is that a "mu" of 5?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=poisson+%285%2C5%2C8%29 i cant seem to get this thing to read my butchering corretctly :/

OpenStudy (amistre64):

ill take a look at it after class, in case noone else has taken a stab at it, good luck :)

OpenStudy (anonymous):

thank you very much

OpenStudy (amistre64):

from: http://stattrek.com/probability-distributions/poisson.aspx P(x; μ) = (e-μ) (μx) / x! the cumulative results are then P(x1;u) + P(x2;u) +P(x3;u) + ... +P(xn;u) in this case we want P(6) + P(7) when u=5 P(6; 5) = (e-5) (30) / 6! P(7; 5) = (e-5) (35) / 7! when I add these together I get -0.1109 how this relates to a table tho, or even if i did it righ ... i dunno :)

OpenStudy (amistre64):

i see an error, when i copy pasted the formula it didnt read the ^ part :) \[P(x;u)=e^{-u}\ \frac{xu}{x!}\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=e%5E%28-5%29%2830%2F6%21%2B35%2F7%21%29 .0003 if that might be useful lol